OpenStudy (anonymous):

sqrt (1-5\9) -sqrt ( 2+1\4) +sqrt (3+1\16) =

OpenStudy (anonymous):

What is \[1 - \frac{ 5 }{ 9 }\] as a single fraction?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

No I'm asking you, what is it?

OpenStudy (anonymous):

is the square root and must perform actions to find the value which should be one of these" 2: 3\2 : 11\12 or 2\3

OpenStudy (anonymous):

No no, I know what the question is asking, I am trying to help you work the answer out. We are looking at the first term: \[\sqrt{1 - \frac{ 5 }{ 9 }}\] we are going to simplify this. Do you know how to simplify \[1 - \frac{ 5 }{ 9 }\] by using a common denominator?